If the consumption function is given by C = 150 + 0.85(Y - T) and T increases by 1 unit, then savings: a. decreases by 0.85 units. b. decreases by 0.15 units. c. increases by 0.15 units. d. increases by 0.85 units.
Answer #1

b. Decreases by 0.15 units

From the consumption function we can tell that MPC= 0.85

MPS= 1-MPC = 1-0.85= 0.15

Change in savings =- MPS*change in T = -0.15*1= -.15